# Rethinking Fine-Tuning when Scaling Test-Time Compute: Limiting Confidence Improves Mathematical Reasoning

### v2

@@ 5.4 Approximate DCO also improves MATH answering with chain-of-thought reasoning
In Section 5.2 on solving problems in MATH, the model is trained to directly give an answer $y$ to a problem $x$. In this case, one can implement DCO at training time by explicitly computing the model confidence $\hat{p}(y|x)$ and using it in Equation 4. However, in an alternate powerful Chain-of-Thought (CoT) training paradigm, the training data consists of triplets $(x,c,y)$ where $x$ and $y$ are the problem and answer as before, but now $c$ is a CoT reasoning trace that explains how to derive $y$ from $x$. The model is trained on the triplet $(x,c,y)$, and at test time, when given a new problem $x^{\prime}$, it generates a reasoning trace $c^{\prime}$ and answer $y^{\prime}$. Importantly, the model is evaluated on whether its answer $y^{\prime}$ is correct independent of the reasoning trace $c^{\prime}$ it emits.
@@
Thus, to estimate the probability $\hat{p}(y|x)$ that the model assigns to any answer $y$, one can no longer compute it directly as in the direct answer case in Section 5.2. Instead, one must marginalize over all possible reasoning traces $c$ to obtain $\hat{p}(y|x)=\sum_{c}\hat{p}(y,c|x)$. Because exact marginalization is intractable, to employ DCO in Equation 4 in the CoT setting, we replace the marginalization with a Monte Carlo estimate of $\hat{p}(y|x)$, and insert this estimate into the overconfidence regularization factor $F$ in Equation 4. We call this algorithm approximate DCO, denoted by DCOa. We also compute Monte Carlo estimates of the probability the model assigns to its most likely final answer, $\hat{p}(y^{\mathrm{mode}}|x)$, as a proxy for the greedy completion probability $\hat{p}(y^{\mathrm{greedy}}|x)$ used to study overconfidence in the direct answer setting in  Section 5.2. Note that in the CoT setting the greedy answer would correspond to a single most likely reasoning trace, but we want the probability of the most likely final answer obtained by marginalizing over all reasoning traces.
@@
We conduct experiments on the MATH dataset. For the baseline experiment, we first fine-tune a Llama-3-8B-base model with CE loss on the golden CoT solutions. We find in Table 3 that pass@1 coverage robustly improves over the course of CE loss training. Intriguingly, in the CoT setting, overfitting induced by the misalignment of CE loss at training time with the pass@N strategy at test time, is not as severe as in the direct answer setting in Table 1. If overfitting indeed arises from the model overconfidence induced by CE loss, we would expect that minimizing CE loss in the CoT setting does not make the model as overconfident as it does in the direct answer case. This is indeed confirmed: compare Figure 6 left, in the CoT case, where the confidence of the modal answer shifts slightly to the right over epochs, to Figure 1 leftmost, in the direct answer case, where it shifts much further right.
We next explore the properties of DCOa with $N=64$. We find that training with DCOa ($N^{\prime}=64$) underperforms relative to training with CE loss for pass@1 at test time, but outperforms when the test strategy is pass@N for larger $N$ (Table 3). This is similar to the direct answer behavior in Figure 4. The largest improvement occurs when training with DCOa at $N^{\prime}=64$ is aligned with pass@N at $N=64$. Furthermore, our theory in Section 4.2 predicts that DCOa training should lower model confidence relative to CE loss training, thereby explaining the improved pass@64 performance. We confirm this prediction in Figure 6 right, which reveals that DCOa successfully prevents the model from becoming overconfident in its most likely answer $y^{\mathrm{mode}}$.
@@
Table 3: Pass@N test coverage on MATH obtained from fine-tuning Llama-3-8B-base on CoT traces using CE or DCOa loss with $N=64$. $\pm$ error is standard error of the mean.
pass@1 pass@16 pass@64
Epoch 1 CE 4.3 $\pm$ 0.1% 30.2 $\pm$ 0.7% 51.2 $\pm$ 1.2%
DCOa 5.0 $\pm$ 0.1% 34.2 $\pm$ 0.6% 55.3 $\pm$ 1.0%
Epoch 3 CE 9.0 $\pm$ 0.1% 41.6 $\pm$ 0.4% 61.7 $\pm$ 0.4%
DCOa 8.3 $\pm$ 0.1% 42.6 $\pm$ 0.6% 63.1 $\pm$ 0.8%
Epoch 5 CE 9.9 $\pm$ 0.1% 40.9 $\pm$ 0.4% 60.9 $\pm$ 0.6%
DCOa 7.8 $\pm$ 0.1% 42.8 $\pm$ 0.5% 64.3 $\pm$ 0.6%
Figure 6: Overconfidence persists in CoT fine-tuning with CE loss, which DCOa successfully limits. We fine-tune a Llama-3-8B base model on the MATH dataset on CoT traces. We plot the distribution of the estimated model confidences $\hat{p}(y^{\mathrm{mode}}|x)$ over samples in the test set at various points in training. Left: for CE loss, the model becomes more confident in its most likely answers, as the confidence distribution shifts slightly to the right as training progresses. This effect is milder than in the direct answer setting (Figure 1). Right: for DCOa, the rightward shift in the confidence distribution of the most likely answer over training time is limited; DCOa again limits overconfidence relative to CE loss.
@@ 6 Discussion
In summary, all our results suggest the need for a tight co-design of two traditionally separate phases of LLM development: (1) model training or fine-tuning and (2) test-time search/reasoning strategies and budget. If the former is misaligned with the latter, then more training can actually impair performance gains from scaling test-time compute. But if they are properly co-designed, end-to-end training and test time performance can be far more effective. We have shown how to modify standard cross-entropy loss for training to be better aligned to a pass@N strategy for large N at testing. Moreover, we have suggested and empirically confirmed why this co-design of training loss and pass@N strategy is essential, because optimal policies for pass@N at large N should unconfidently explore while the optimal policies for pass@N at small N should confidently exploit.
This notion of co-design opens up many more theoretical and empirical questions associated with different test-time strategies such as self-verification and MCTS. Furthermore, one can go beyond test-time search to recursive self-improvement, where new solutions found by search are filtered and then used to retrain the model. A theoretical understanding of the capabilities achievable by recursive self-improvement through co-design of training, search, and filtering, remains an outstanding research problem.
Acknowledgments
@@ A.1 Proof of approximately well calibration under optimal policy
Second, given $p_{1}$ and $\sum_{i=1}^{k}p_{i}\geq 1-\epsilon$, we have
$\pi_{1}\left(\underset{\{\hat{p}_{i}\}_{i=1}^{k}}{\text{argmin}}\{\sum_{i=1}^{k}p_{i}(1-\hat{p}_{i})^{N}+\sum_{i=k+1}^{R}p_{i}\}\right)\leq\pi_{1}\left(\underset{\{\hat{p}_{i}\}_{i=1}^{k}}{\text{argmin}}\{p_{1}(1-\hat{p}_{1})^{N}+\sum_{i=2}^{k}\frac{1-p_{1}-\epsilon}{k-1}(1-\hat{p}_{i})^{N}+\sum_{i=k+1}^{R}p_{i}\}\right).$ (11)
@@
$\hat{p}_{1}^{*}\leq\underset{\hat{p}_{1}}{\text{argmin}}\{p_{1}(1-\hat{p}_{1})^{N}+(1-p_{1}-\epsilon)(1-\frac{1-\hat{p}_{1}}{k-1})^{N}\}=1-\frac{k-1}{(k-1)^{\frac{N}{N-1}}p_{1}^{\frac{1}{N-1}}\left(1-p_{1}-\epsilon\right){}^{\frac{1}{1-N}}+1}.$ (14)
@@ Appendix B Experimental details
All experiments with model size smaller than 10B are performed on machines with 8 NVIDIA H100 GPUs or 8 NVIDIA A100 GPUs. Experiments with model size larger than 10B are performed on machines with 8 NVIDIA H200 GPUs or 8 NVIDIA B200 GPUs. Our codebase uses PyTorch (Ansel et al., 2024), Accelerate (Gugger et al., 2022), and deepspeed (https://github.com/microsoft/DeepSpeed) to enable efficient training with memory constraints, and vllm (Kwon et al., 2023) for efficient inference. Code will be made available here.
MATH. For experiments with MATH dataset, we fine-tune the Llama-3-8B-base (Grattafiori et al., 2024) on the MATH (Hendrycks et al., 2021) dataset. We start from the base model rather than Llama-3-8B-Instruct to avoid potential leakage of the MATH dataset into Llama-3-8B-Instruct through post-training process. We follow Lightman et al. (2024) and use $12,000$ problems for training and the remaining $500$ for testing. In Sections 4 and 5.2 and Figures 1 and 4, we fine-tune the model for 4 epochs with a learning rate of 2e-5 and batch size 64. We adopt a linear learning rate warmup in the first 20 steps. For experiments with DCO, some of the data may have an extreme confidence regularizer value, if the model is already quite confident in answers to certain problems. To maintain an approximately fixed batch size, we set a threshold of $0.3$ on the confidence regularizer $F$. Any training data examples with an $F$ lower than this threshold will be replaced with new training examples to construct the batch. For the CoT experiments in Section 5.4, we use learning rate 2e-5 and batch size 128, with the same learning rate warmup.
Theorem proving. We adopt the random train and test split as introduced in Yang et al. (2023). The random test set includes 2,000 theorems. We fine-tune the model Qwen2.5-Math-1.5B (Yang et al., 2024) on the training set for 3 epochs with learning rate 1e-5 and batch size 64. We adopt a linear learning rate warmup in the first 20 steps. To evaluate the model, we use LeanDojo (Yang et al., 2023) to interact with the proof assistant. We impose a maximum wall clock time for each theorem, in addition to limiting the number of passes per problem. For experiments with 4k passes, the time budget is fixed at 5,000 seconds. In order to avoid the model going infinitely deep in the search tree, we limit the number of proof steps to be at most 50.
@@
Implementing online inference. Throughout our experiments and analysis, we extensively use the open-source vllm package (Kwon et al., 2023) for efficient inference. Integrating inference into the training loop to enable training under the DCOa objective, as in Section 5.4, poses a challenging implementation problem. We solve this problem by placing vllm worker processes, which together perform inference on all GPUs, in a separate process group and use Ray (https://github.com/ray-project/ray) to isolate them from the training loop. This enables running concurrent training and inference on the same set of GPUs. We believe this inference in the loop setup will be useful to the community, as it enables straightforward implementation of online data filtering approaches for LLM training.
Computational cost. For both DCO and DCOstep, the additional computational cost is minimal—limited to computing the DCO factor, which is negligible compared to the cost of forward and backward passes.
However, DCOa, when applied to fine-tuning with CoT traces, introduces significant overhead due to the need for online Monte Carlo estimation. In the setting of Llama-3 8B, with 32 layers, hidden dimension 4096, context length 1024, and batch size 128, we estimate 5 petaflops for each forward-backward pass. Generating $N_{MC}=64$ CoT’s for each of 128 examples, in turn, requires approximately 50 petaflops. The actual inference compute cost depends on model performance: stronger models, which will discard more samples, will need to process more samples in order to construct a full batch for the next training step. Therefore there is a tradeoff: constructing a batch of 128 hard samples requires running inference on more samples, but also means fewer steps are required to make a full pass through the dataset.
@@ Appendix C Additional results
C.1 Optimality of DCO is achieved when the training objective aligns with the test objective.

### v3

@@ 5.4 Approximate DCO also improves performance with chain-of-thought reasoning
In Section 5.2, the model is trained to directly give an answer $y$ to a problem $x$. In this case, one can implement DCO at training time by explicitly computing the model confidence $\hat{p}(y|x)$ and using it in Equation 4. However, in a Chain-of-Thought (CoT) training paradigm, the training data consists of triplets $(x,c,y)$ where $x$ and $y$ are the problem and answer as before, but now $c$ is a CoT reasoning trace that explains how to derive $y$ from $x$. The model is trained on the triplet $(x,c,y)$, and at test time, given a new problem $x^{\prime}$, it generates a reasoning trace $c^{\prime}$ followed by an answer $y^{\prime}$. Importantly, the model is evaluated on whether its answer $y^{\prime}$ is correct independent of its reasoning trace $c^{\prime}$.
@@
Thus, to estimate the probability $\hat{p}(y|x)$ that the model assigns to any answer $y$, one can no longer compute it directly as in the direct answer case. Instead, one must marginalize over all possible reasoning traces $c$ to obtain $\hat{p}(y|x)=\sum_{c}\hat{p}(y,c|x)$. Exact marginalization is intractable, so we replace the marginalization with a Monte Carlo estimate of $\hat{p}(y|x)$, which we insert into the overconfidence regularization factor $F$ in Equation 4. We call this algorithm approximate DCO, denoted by DCOa. We also compute Monte Carlo estimates of the probability the model assigns to its most likely final answer, $\hat{p}(y^{\mathrm{mode}}|x)$, as a proxy for the model’s max confidence. Note that in the CoT setting the greedy answer would correspond to a single most likely reasoning trace, but we want the probability of the most likely final answer marginalizing over all reasoning traces.
@@
We conduct experiments on the MATH dataset. As a baseline, we first fine-tune a Llama-3-8B-base model using CE loss on the golden CoT solutions. We find in Table 3 that pass@1 coverage robustly improves over the course of CE loss training. Intriguingly, in the CoT setting, overfitting induced by the misalignment of CE loss at training time with the pass@N strategy at test time, appears less severe compare with the direct answer setting (cf. Table 1). Nevertheless, we still find a clear increase in the model’s max confidence through training (Figure 3 (a)). In contrast, training with DCOa mitigates the rise in the model’s max confidence (Figure 3 (b)). As with the direct answer setup, model trained with DCOa ($N^{\prime}=64$) exhibits lower pass@1 performance compared to the CE baseline but demonstrates superior performance when evaluated with pass@N for larger $N$ values (Table 3).
To provide a comprehensive comparison, we also evaluate a model trained with reinforcement learning. Specifically, we use GRPO [27] to train the base model with the r1 [10] prompt template. Interestingly, we observe that the GRPO-trained model exhibits pronounced overconfidence (Figure 3 (c)), yet only a small portion of the model’s highly confident answers are correct (Figure 3 (d)). This negatively impacts the model’s pass@N performance when scaling up $N$ (Table 3). Although GRPO demonstrates superior pass@1 performance, the model trained with DCOa achieves a significantly stronger performance when evaluated with pass@N for larger $N$ values (Table 3).
@@
Table 3: Pass@N coverage on MATH ($\pm$ indicates standard error of the mean). We fine-tune Llama-3-8B-base on CoT traces. Training with DCOa using $N^{\prime}=64$ underperforms CE loss for pass@1, but outperforms it for higher $N$, with the largest improvement when $N^{\prime}=N$. The GRPO model has the strongest pass@1 performance, but its performance degrades at larger $N$ due to overconfidence.
Epoch 1 Epoch 3 Epoch 5
Pass@N CE GRPO DCOa CE GRPO DCOa CE GRPO DCOa
Pass@1 $4.3^{\pm 0.1}$ 3.0±0.1 5.0±0.1 9.0±0.1 12.6±0.6 8.3±0.1 9.9±0.1 14.6±0.1 7.8±0.1
Pass@16 30.2±0.7 27.1±0.7 34.2±0.6 41.6±0.4 35.8±0.3 42.6±0.6 40.9±0.4 29.2±0.2 42.8±0.5
Pass@64 51.2±1.2 50.6±0.7 55.3±1.0 61.7±0.4 46.3±0.4 63.1±0.8 60.9±0.6 35.4±0.3 64.3±0.6
@@ 6 Discussion
In summary, all our results suggest the need for a tight co-design of two traditionally separate phases of LLM development: (1) model training or fine-tuning and (2) test-time search/reasoning strategies and budget. If the former is misaligned with the latter, then more training can actually impair performance gains from scaling test-time compute. But if they are properly co-designed, end-to-end training and test time performance can be far more effective. We have shown how to modify standard cross-entropy loss for training to be better aligned to a pass@N strategy for large $N$ at test-time. Moreover, we have suggested and empirically confirmed why this co-design of training loss and pass@N strategy is essential, because optimal policies for pass@N at large $N$ should unconfidently explore while the optimal policies for pass@N at small $N$ should confidently exploit.
Limitations. Our study primarily investigates SFT with the pass@N test-time strategy for its simplicity. However, this strategy requires a verifier, restricting its applicability to problems where verification is feasible. Extending our findings to other test-time strategies, such as best-of-N, is an important direction for future research. Additionally, while we have noted similar overconfidence phenomena in GRPO, a more comprehensive study of RL methods is beyond the current scope. Lastly, our work only focuses on mathematical reasoning; generalizing our findings to other domains, such as coding tasks where unit tests serve as natural verifiers, is left for future work.
Acknowledgments and Disclosure of Funding
@@ A.1 Proof of approximately well calibration under optimal policy
Second, given $p_{1}$ and $\sum_{i=1}^{k}p_{i}\geq 1-\epsilon$, $\pi_{1}\left(\underset{\{\hat{p}_{i}\}_{i=1}^{k}}{\text{argmin}}\{\sum_{i=1}^{k}p_{i}(1-\hat{p}_{i})^{N}+\sum_{i=k+1}^{R}p_{i}\}\right)$ is bounded from above by
$\pi_{1}\left(\underset{\{\hat{p}_{i}\}_{i=1}^{k}}{\text{argmin}}\{p_{1}(1-\hat{p}_{1})^{N}+\sum_{i=2}^{k}\frac{1-p_{1}-\epsilon}{k-1}(1-\hat{p}_{i})^{N}+\sum_{i=k+1}^{R}p_{i}\}\right).$ (11)
@@
$\hat{p}_{1}^{*}&\leq\underset{\hat{p}_{1}}{\text{argmin}}\{p_{1}(1-\hat{p}_{1})^{N}+(1-p_{1}-\epsilon)(1-\frac{1-\hat{p}_{1}}{k-1})^{N}\}\\ &=1-\frac{k-1}{(k-1)^{\frac{N}{N-1}}p_{1}^{\frac{1}{N-1}}\left(1-p_{1}-\epsilon\right){}^{\frac{1}{1-N}}+1}\end{aligned}.$ (14)
@@ Appendix B Experimental details
Our codebase uses PyTorch [45], Accelerate [46], and deepspeed (https://github.com/microsoft/DeepSpeed) to enable efficient training with memory constraints, and vllm [47] for efficient inference. Code will be made available at https://github.com/allanraventos/refine.
MATH. For experiments with MATH dataset [41] (license: MIT), we fine-tune the Llama-3-8B-base [40] on the MATH [41] dataset. We start from the base model rather than Llama-3-8B-Instruct to avoid potential leakage of the MATH dataset into Llama-3-8B-Instruct through post-training process. We follow Lightman et al. [22] and use $12,000$ problems for training and the remaining $500$ for testing. In Sections 4 and 5.2, Figure 1, Figure 2 (a) and (b), we fine-tune the model for 4 epochs with a learning rate of 2e-5 and batch size 64. We adopt a linear learning rate warmup in the first 20 steps. For experiments with DCO, some of the data may have an extreme confidence regularizer value, if the model is already quite confident in answers to certain problems. To maintain an approximately fixed batch size, we set a threshold of $0.3$ on the confidence regularizer $F$. Any training data examples with an $F$ lower than this threshold will be replaced with new training examples to construct the batch. For the CoT experiments in Section 5.4, we use learning rate 2e-5 and batch size 128, with the same learning rate warmup.
Theorem proving. We use LeanDojo benchmark[42] (license: CC-BY 2.0) and adopt the random train and test split as introduced in Yang et al. [42]. The random test set includes 2,000 theorems. We fine-tune the model Qwen2.5-Math-1.5B [26] on the training set for 3 epochs with learning rate 1e-5 and batch size 64. We adopt a linear learning rate warmup in the first 20 steps. To evaluate the model, we use LeanDojo [42] to interact with the proof assistant. We impose a maximum wall clock time for each theorem, in addition to limiting the number of passes per problem. For experiments with 4k passes, the time budget is fixed at 5,000 seconds. In order to avoid the model going infinitely deep in the search tree, we limit the number of proof steps to be at most 50.
@@
GRPO. We perform GRPO fine-tuning [10, 27] on problems of all difficulty levels in the MATH training set. We follow the modifications to GRPO proposed in Dr. GRPO ([31]), that is, we do not normalize advantages by the standard deviation of the group rewards, do not normalize a roll-out’s contribution to the loss by its length, and do not include a KL divergence term with respect to the reference model. Furthermore, we perform only one policy gradient update for each set of model roll-outs, so there is no need for clipping. This leads to the surrogate objective taking the form,
$\mathcal{L}(\theta)=\frac{1}{N^{\prime}}\sum_{i=1}^{N^{\prime}}\sum_{t=1}^{|o_{i}|}\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\mathrm{stop\_grad}\left(\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\right)}\hat{A}_{i,t}$
where we adopt notation similar to [31]; $N^{\prime}$ is the number of roll-outs for a single problem statement $q$, $o_{i}$ is the $i$th roll-out in the group, $o_{i,t}$ is the $t$th token in $o_{i}$, and $\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})$ is the probability assigned by the model to token $o_{i,t}$ conditioned on context $o_{i,<t}$. The $\mathrm{stop\_grad}$ operator prevents the flow of gradients when computing $\nabla_{\theta}\mathcal{L}$. The advantages, in turn, are computed as $\hat{A}_{i,t}=r(q,o_{i})-\frac{1}{N^{\prime}}\sum_{i^{\prime}=1}^{N^{\prime}}r(q,o_{i^{\prime}})$, where $r(q,o_{i})$ is the binary reward assigned by the verifier to the solution $o_{i}$ for the problem $q$.
For our experiments, we use $N^{\prime}=8$ and batch size 128, which results in a total of 1,024 roll-outs per batch. We also perform batch balancing such that each batch contains no problems for which either $r(q,o_{i})=0$ for all $i$ or $r(q,o_{i})=1$ for all $i$, since these samples do not provide any gradient signal. We use a constant learning rate of 1e-6,with a 20-step linear warmup; we train for 5 epochs. Note that batch balancing results in fewer gradient steps. Additionally, we use the r1 template from [10] for our experiments,
A conversation between User and Assistant. The User asks a question, and the Assistant solves it. The Assistant first thinks about the reasoning process in the mind and then provides the User with the answer. The reasoning process is enclosed within <think> </think> and the answer is enclosed within <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think> <answer> answer here </answer>.
User: {question}
Assistant: <think>
We strictly require solutions to follow the format specified in the prompt.
@@
Implementing online inference. Throughout our experiments and analysis, we extensively use the open-source vllm package [47] for efficient inference. Integrating inference into the training loop to enable training under the DCOa objective, as in Section 5.4, poses a challenging implementation problem. We solve this problem by placing vllm worker processes, which together perform inference on all GPUs, in a separate process group and use Ray (https://github.com/ray-project/ray) to isolate them from the training loop. This enables running concurrent training and inference on the same set of GPUs. We believe this inference in the loop setup will be useful to the community, as it enables straightforward implementation of online data filtering approaches for LLM training.
Computational cost. All experiments with model size smaller than 10B are performed on machines with 8 NVIDIA H100 GPUs or 8 NVIDIA A100 GPUs. Experiments with model size larger than 10B are performed on machines with 8 NVIDIA H200 GPUs or 8 NVIDIA B200 GPUs. Specifically, for DCO on the MATH dataset, each run required approximately 1 NVIDIA H100 GPU-hour when using the Llama-3-8B-base model, and approximately 4 NVIDIA B200 GPU-hours when using the Llama-3-70B-base model. For DCOstep on the LeanDojo benchmark with the Qwen2.5-Math-1.5B model, each run required around 12 NVIDIA H100 GPU-hours. For DCOa on the MATH dataset with CoT, each run consumed about 60 NVIDIA H100 GPU-hours for the Llama-3-8B-base model.
For both DCO and DCOstep, the additional computational cost is minimal—limited to computing the DCO factor, which is negligible compared to the cost of forward and backward passes. However, DCOa, when applied to fine-tuning with CoT traces, introduces significant overhead due to the need for online Monte Carlo estimation. In the setting of Llama-3 8B, with 32 layers, hidden dimension 4096, context length 1024, and batch size 128, we estimate 5 petaflops for each forward-backward pass. Generating $N_{MC}=64$ CoT’s for each of 128 examples, in turn, requires approximately 50 petaflops. The actual inference compute cost depends on model performance: stronger models, which will discard more samples, will need to process more samples in order to construct a full batch for the next training step. Therefore there is a tradeoff: constructing a batch of 128 hard samples requires running inference on more samples, but also means fewer steps are required to make a full pass through the dataset.
@@ Appendix C Additional results
C.1 Easy data drives overconfidence
We have shown in Section 4.2 that overconfidence limits performance gains from scaling test-time compute. Here we investigate whether all training data contribute equally to the drop in pass@N test performance. To quantify this effect, we define the test loss as the negative log-probability of coverage, $\mathcal{L}_{\text{test}}^{N}=-\sum_{(x,y)\in\mathcal{D}^{\mathrm{test}}}\mathcal{\log}\ \mathcal{C}_{(x,y)}^{N}$, where $\mathcal{C}_{(x,y)}^{N}$ denotes the coverage for a single test example $(x,y)$. To measure how a training batch $\mathcal{B}$ influences the final test loss, we compute the directional derivative $\nabla_{g}\mathcal{L}_{\text{test}}^{N}$ for the gradient direction $g=-\sum_{(x_{i},y_{i})\in\mathcal{B}}\nabla\ell_{\text{CE}}(x_{i},y_{i})$, where $\ell_{\mathrm{CE}}$ is the standard CE loss on a single example. A negative value, $\nabla_{g}\mathcal{L}_{\text{test}}^{N}<0$, indicates that a gradient step in the direction of $g$ decreases the test loss.
C.2 Optimality of DCO is achieved when the training objective aligns with the test objective.